Kelly’s Jewelry has the following transactions during the year: total jewelry sales = $680,000; sales discounts = $16,500; sales returns = $43,000; sales allowances = $23,000. In addition, at the end of the year the company estimates the following transactions associated with jewelry sales in the current year will occur next year: sales discounts = $1,650; sales returns = $5,160; sales allowances = $3,070. Compute net sales.
Answer #1
Total sales 680000
Less: Sales discounts (18150) =16500+1650
Less: Sales return (48160) =43000+5160
Less: Sales allowances (26070) =23000+3070
Net sales 587620